By how much should the temperature of an aluminum sample
change if its volume is to change by 1.5 % of its volume.
(Coefficient of linear expansion of aluminum is 2.4e-5 ⁄ C)


The answer is: 208.333 C but I do not know how to get there without having the volume or radius.
1
Expert's answer
2019-07-09T13:08:22-0400


In this task, you must assume that we have an aluminum cube volume 1 m3.

The dependence of the volume on the length of the edge of the cube is calculated by the formula:


"V = S^3"

"S = \sqrt[3]{V}"

"S = 1 m"


Cube volume after heating:


"V_1 = V + V*0.015"


"V_1 = 1 + 0.015 = 1.015m^3"




"S_1 = \sqrt[3]{1.015}m"


"S_1 = 1.005 m"



Finally:

"\alpha = {\Delta S \over S\Delta T}"


"\Delta S = S_1 - S"

"\Delta S = 1.005 - 1 = 0.005 = 5*10^{-3}"


"\Delta T = {\Delta S \over \alpha S }"

"\Delta T = {5*10^{-3} \over 2.4*10^{-6}} = {0.005 \over 0.000024} = 208.333 K"


Answer: 208.333 K